# Influence of thermal noise on the field-driven dynamics of the non-collinear antiferromagnet Mn3Sn

### v1

@@ Abstract
Non-collinear chiral antiferromagnets (AFMs) of the form Mn3X (X = Sn, Ge, Ir, and Pt) Shukla, Qian, and Rakheja (2025); Rimmler, Pal, and Parkin (2025) have gained significant attention owing to their large anomalous Hall effect (AHE), Nakatsuji, Kiyohara, and Higo (2015); Nayak et al. (2016); Chen, Niu, and MacDonald (2014) spin Hall effect (SHE), Zhang et al. (2016) anomalous Nernst effect (ANE), Ikhlas et al. (2017); Hong et al. (2020) magneto-optical Kerr effect (MOKE), Higo et al. (2018) and finite tunneling magnetoresistance (TMR). Dong et al. (2022); Chen et al. (2023); Qin et al. (2023); Wang et al. (2024) The anomalous Hall conductivity (AHC) of Mn3Sn, a prototypical non-collinear chiral AFM, can range from 30 to 40  $\Omega^{-1}\cdot$cm-1 at room temperature, which is attributed to the breaking of time-reversal symmetry (TRS). Liu et al. (2023); Yano et al. (2024) In its bulk form and below its Néel temperature ($T_{\mathrm{N}}$) of approximately 420 K, Mn3Sn exhibits a six-fold degenerate energy landscape. Higo et al. (2018); Sung et al. (2018) However, in an epitaxially grown Mn3Sn/MgO(110) [001] bilayer structure, the energy landscape is found to possess two-fold symmetry with the emergence of a perpendicular magnetic anisotropy (PMA) owing to an epitaxial tensile strain. Higo et al. (2022); Yoon et al. (2023)
@@
Current-driven dynamics, including deterministic switching and chiral oscillations, have been experimentally explored in polycrystalline as well as epitaxial films of thicknesses ranging from sub-10 nm to 100 nm. Tsai et al. (2020); Takeuchi et al. (2021); Yan et al. (2022); Pal et al. (2022); Krishnaswamy et al. (2022); Higo et al. (2022); Yoon et al. (2023); Xu et al. (2023); Yoo et al. (2024); Zheng et al. (2025); Lee et al. (2025) The AHE signal in the case of polycrystalline films, measured in a typical Hall bar setup, is considered to be an average of the final magnetic states in the various composing grains. Ikeda et al. (2018); Xie et al. (2022) While in some cases pure spin-orbit torque (SOT) is considered to be the fundamental mechanism behind the observed dynamics, Tsai et al. (2020); Takeuchi et al. (2021); Higo et al. (2022); Yoon et al. (2023) others have attributed the observed changes in the AHE signal to a combination of Joule heating and SOT (i.e., demagnetization of the polycrystalline Mn3Sn film as the temperature rises above $T_{\mathrm{N}}$ due to Joule heating, followed by SOT-driven re-magnetization as the film cools below $T_{\mathrm{N}}$ during the slow decrease in current. Pal et al. (2022); Krishnaswamy et al. (2022); Yoo et al. (2024) The SOT-driven oscillation Zhao, He, and Cai (2021); Shukla and Rakheja (2022); Lund et al. (2023) and switching Shukla, Qian, and Rakheja (2023); Xu et al. (2024); Shukla, Qian, and Rakheja (2024); Zhao et al. (2025) dynamics in both six-fold and two-fold energy degenerate Mn3Sn crystals have also been explored theoretically. In particular, these works have extended the ferromagnetic theory to shed light on the threshold current requirement for initiating dynamics as well as the dependence of the various threshold currents on the material properties. They have also explored the possible dependence of the oscillation frequency and the switching time on the input stimuli and the material parameters. The manipulation of the order parameter in Mn3Sn via SOT, together with a finite TMR in all-Mn3Sn junctions, can enable the development of AFM-based spintronic devices that are fully compatible with electronic circuitry.
@@
Figure 1: (a) Kagome lattice of Mn3Sn. (b) Octupole moment and relevant angles in Cartesian coordinate system.
@@
In this paper, we investigate the impact of thermal noise on the stability of a strained $\mathrm{Mn_{3}Sn}$ bit in the presence of an external magnetic field. We present a comprehensive analysis of the mechanisms underlying the free-energy symmetry breaking, examining the dependence of free energy on both the magnitude and orientation of the applied field, as well as the associated fluctuation modes. The interwell escape time of the magnetic octupole moment is theoretically evaluated as a function of the external field and temperature, with results corroborated by numerical simulations in the low-energy barrier regime ($\Delta E_{0}\leq 5k_{\mathrm{B}}T$). We further extend our theoretical framework to encompass the relaxation dynamics of octupole moment and discuss prospective applications in random number generation and probabilistic computing.
Mn3Sn crystallizes into a hexagonal Kagome $D0_{19}$ lattice below its Néel temperature, exhibiting a slight excess of Mn atoms. Shukla, Qian, and Rakheja (2025) Sn atoms occupy the centers of the hexagons, whereas Mn atoms are located at the vertices. Within each Kagome layer, the Mn atoms contribute three magnetic moments that align at $120^{\circ}$ angles, forming a geometrically frustrated, noncollinear triangular spin structure. These moments are slightly canted toward the in-plane easy axis, giving rise to a non-zero net magnetization as shown in Fig. 1(a). The magnetic ground state typically exhibits six-fold degeneracy; Shukla, Qian, and Rakheja (2023) however, in the presence of an in-plane uniaxial strain, the symmetry is reduced to two-fold. Higo et al. (2022); Yoon et al. (2023)
@@
$\mathcal{H}(\mathbf{m})$ $=J_{\mathrm{E}}((1+\delta_{\mathrm{E}})\mathbf{m}_{1}\cdot\mathbf{m}_{2}+\mathbf{m}_{2}\cdot\mathbf{m}_{3}+\mathbf{m}_{3}\cdot\mathbf{m}_{1})+D_{\mathrm{M}}\mathbf{z}\cdot(\mathbf{m}_{1}\times\mathbf{m}_{2}+\mathbf{m}_{2}\times\mathbf{m}_{3}+\mathbf{m}_{3}\times\mathbf{m}_{1})-\sum_{i=1}^{3}{\left[K_{\mathrm{u}}(\mathbf{m}_{i}\cdot\mathbf{u}_{i})^{2}+M_{\mathrm{s}}\mathbf{H}_{\mathrm{a}}\cdot\mathbf{m}_{i})\right]},$ (1)
@@
$\mathcal{H}(\theta_{\mathrm{oct}},\varphi_{\mathrm{oct}})=\frac{3}{4}M_{\mathrm{s}}H_{\mathrm{K}}\cos(2\varphi_{\mathrm{oct}})+\frac{3}{2}M_{\mathrm{s}}H_{\mathrm{J}}\cos^{2}\theta_{\mathrm{oct}}-\frac{M_{\mathrm{s}}H_{\mathrm{a}}\cos{(\theta_{\mathrm{H}})}}{J_{\mathrm{E}}+\sqrt{3}D_{\mathrm{M}}}\Big{[}K_{\mathrm{u}}\cos{(\varphi_{\mathrm{oct}}-\varphi_{\mathrm{H}})}+J_{\mathrm{E}}\delta_{E}\cos{(\varphi_{\mathrm{oct}}+\varphi_{\mathrm{H}})}\Big{]}.$ (2)
Here, we used $\mathbf{m}_{\mathrm{oct}}=\frac{1}{3}\mathcal{M}_{zx}\quantity[R\left(\frac{2\pi}{3}\right)\mathbf{m}_{1}+R\left(-\frac{2\pi}{3}\right)\mathbf{m}_{2}+\mathbf{m}_{3}]$ while $H_{\mathrm{J}}=\frac{(3J_{\mathrm{E}}+\sqrt{3}D_{\mathrm{M}})}{M_{s}}$ and $H_{\mathrm{K}}=-\frac{(4K_{\mathrm{u}}J_{\mathrm{E}}\delta_{\mathrm{E}})}{3M_{s}(J_{\mathrm{E}}+\sqrt{3}D_{\mathrm{M}})}$ are the strength of exchange field and uniaxial anisotropy field, respectively. $\theta_{\mathrm{H}}$ and $\varphi_{\mathrm{H}}$ are the polar and the azimuthal angle of the applied external field $\mathbf{H}_{\mathrm{a}}$, respectively.
@@
$\delta_{\mathrm{E}}$ Strain parameter $-2.6\times 10^{-4}$
@@
From Eq. (2), strong exchange interactions constrain $\mathbf{m}_{\mathrm{oct}}$ to lie within the Kagome plane, resulting in two degenerate energy maxima at $\varphi_{\mathrm{oct}}=0$ and $\pi$, and two minima at $\varphi_{\mathrm{oct}}=\pi/2$ and $23\pi/$ 2, in the absence of an external field. The equilibrium states near $\varphi_{\mathrm{oct}}=\pi/2$ and $23\pi/$ 2 are hereafter referred to as the up and down states, respectively. An applied external field breaks the symmetry of the local extrema. As shown in Fig. 2(a), when $\mathbf{H}_{\mathrm{a}}$ is applied along the $-x$ direction (i.e., $\varphi_{\mathrm{H}}=\pi$), the degeneracy of the energy maxima is lifted. For a film of volume $\mathcal{V}$, the energy barriers, $\Delta E_{\mathrm{x},1}$ and $\Delta E_{\mathrm{x},2}$, evaluated at $\varphi_{\mathrm{oct}}=0$ and $\pi$, respectively, are given as
@@
where $H_{\mathrm{x,eff}}=H_{\mathrm{ax}}\frac{(K_{\mathrm{u}}+J_{\mathrm{E}}\delta_{\mathrm{E}})}{3(J_{\mathrm{E}}+\sqrt{3}D_{\mathrm{M}})}$ is the normalized external field applied along the $-x$ direction. In this configuration, the up and down states remain energetically equivalent; however, thermal fluctuations preferentially promote transitions across $\Delta E_{\mathrm{x,2}}$ due to its lower energy barrier.
In contrast, when the external field is applied along the $-y$ direction, it primarily perturbs the energy minima, as shown in Fig. 2(b). The energy maxima remain degenerate, resulting in symmetric energy barriers on either side of each state. The analytical expressions for the energy barriers $\Delta E_{\mathrm{y,1}}$ and $\Delta E_{\mathrm{y,2}}$, corresponding to the up and down states, respectively, are
@@
where $H_{\mathrm{y,eff}}=H_{\mathrm{ay}}\frac{(K_{\mathrm{u}}-J_{\mathrm{E}}\delta_{\mathrm{E}})}{3(J_{\mathrm{E}}+\sqrt{3}D_{\mathrm{M}})}$ denotes the normalized external field applied along the $-y$ direction. For comparison, Fig. 2(c) quantitatively illustrates the quadratic dependence of the energy barrier density on the magnitude of the external field. Additionally, we define $\Delta E_{0}$ as the energy barrier height in the absence of an external field, which serves as a reference in subsequent analyses.
Based on the harmonic transition-state theory (HTST), Coffey and Kalmykov (2012); Morse and Feshbach (1946) the escape time with intermediate-to-high (IHD) damping is formulated per Eq. (5). In this equation, $\varepsilon_{j,\mathrm{min(sp)}}$ denotes the $j$-th eigenvalue of the Hessian matrix in the harmonic approximation near the energy minimum (saddle point), and $\lambda_{+}$ is the positive eigenvalue of the linearized Laudau-Lifshitz-Gilbert (LLG) equation at the saddle point. $P_{\mathrm{min(sp)}}$ represents the number of Goldstone modes at the minimum (saddle point), and $V_{\mathrm{min(sp)}}$ denotes the corresponding phase-space volume. Additionally, $A(\delta E/k_{\mathrm{B}}T)$ is the depopulation factor that accounts for the coupling strength to the thermal bath, where $\delta E$ is the energy dissipated along an iso-energy contour encompassing the energy maximum. See Eq. (6) for $A(x)$.
In the theoretical analysis of escape time, we employ the angular variables $\theta_{\mathrm{oct}}$ and $\varphi_{\mathrm{oct}}$ to describe the dynamics of the octupole moment $\mathbf{m}_{\mathrm{oct}}$ within a perturbative framework. He and Liu (2024) Further methodological details are provided in the Supplementary Material . The escape time due to thermal fluctuations over a single saddle point, from the up state to the down state under an external field $H_{\mathrm{ax}}$ or $H_{\mathrm{ay}}$ is given in Eq. (7), where $k=1$ and $k=2$ correspond to the magnetic octupole crossing the saddle point near $\varphi_{\mathrm{oct}}=0$ and $\pi$, respectively. The parameters $\beta_{x}$ and $\beta_{y}$ in Eq. (7) are given as $\left((-1)^{k}\frac{H_{\mathrm{x,eff}}}{H_{\mathrm{K}}}-1\right)$ and $\left(3\frac{H_{\mathrm{y,eff}}^{2}}{H_{\mathrm{K}}^{2}}-1\right)$, respectively. Our calculations suggest that for the magnetic fields and energy barriers considered in this work, $\quantity[A(\delta E/k_{\mathrm{B}}T)]^{-1}$ can be approximated to unity with less than 5% error (see Supplementary Material for details).
@@
[…]
@@
Figure 3: The schematic of escape time from up state to down state while numerical extracted data is in markers and analytical predictions are plotted with lines. $\Delta E_{0}$ ranges from $k_{\mathrm{B}}T$ to $5k_{\mathrm{B}}T$ and field is applied along (a) $x$-axis and (b) $y$-axis. (c) The time evolution of octupole moment with external field and $\Delta E_{0}=4k_{\mathrm{B}}T$.
@@
To validate Eqs. (7) and (8), the dynamics of the magnetic octupole is numerically explored via three coupled stochastic LLG equations, Mayergoyz, Bertotti, and Serpico (2009); Go et al. (2022) for $\Delta E_{0}$ ranging from $k_{\mathrm{B}}T$ to $5k_{\mathrm{B}}T$ and $H_{\mathrm{a}}\leq 0.1~{}\mathrm{T}$. For each sublattice, $i$ (= 1,2,3), the LLG equation is given as $\partialderivative{\mathbf{m}_{i}}{t}=-\gamma(\mathbf{m}_{i}\crossproduct\mathbf{H}_{i}^{\mathrm{eff}})+\alpha\quantity(\mathbf{m}_{i}\times\partialderivative{\mathbf{m}_{i}}{t})$, where $\mathbf{H}_{i}^{\mathrm{eff}}=-\frac{1}{M_{\mathrm{s}}}\frac{\partial\mathcal{H}}{\partial\mathbf{m}_{i}}+\mathbf{H}_{\mathrm{thermal},i}$ is the effective magnetic field and includes the thermal noise field, $\mathbf{H}_{\mathrm{thermal},i}=\sqrt{\frac{2\alpha k_{\mathrm{B}}T}{\gamma\mu_{0}M_{\mathrm{s}}^{2}\mathcal{V}\Delta t}}\bm{\xi}^{i}_{t}$. In this work, we assume $\mathbf{H}_{\mathrm{thermal},i}$ to be Gaussian with zero mean as well as uncorrelated in space, time, and between sublattices, such that $\bm{\xi}^{i}_{t}\sim\mathcal{N}(0,1)$Go et al. (2022); Ament et al. (2016) Here, $\gamma$ and $\mu_{0}$ denote the gyromagnetic ratio and vacuum permeability, respectively, while $\Delta t$ is the time step used in the simulation. Other parameters are defined in Table 1.

### v2

@@ Abstract
Non-collinear chiral antiferromagnets (AFMs) of the form Mn3X (X = Sn, Ge, Ir, and Pt) Shukla, Qian, and Rakheja (2025); Rimmler, Pal, and Parkin (2025) have gained significant attention owing to their large anomalous Hall effect (AHE), Nakatsuji, Kiyohara, and Higo (2015); Nayak et al. (2016); Chen, Niu, and MacDonald (2014) spin Hall effect (SHE), Zhang et al. (2016) anomalous Nernst effect (ANE), Ikhlas et al. (2017); Hong et al. (2020) magneto-optical Kerr effect (MOKE), Higo et al. (2018) and finite tunneling magnetoresistance (TMR). Dong et al. (2022); Chen et al. (2023); Qin et al. (2023); Wang et al. (2024) The anomalous Hall conductivity (AHC) of Mn3Sn, a prototypical non-collinear chiral AFM, can range from 30 to 40  $\Omega^{-1}\cdotcm$-1 at room temperature, which is attributed to the breaking of time-reversal symmetry (TRS). Liu et al. (2023a); Yano et al. (2024) In its bulk form and below its Néel temperature ($T_{\mathrm{N}}$) of approximately 420 K, Mn3Sn exhibits a six-fold degenerate energy landscape. Higo et al. (2018); Sung et al. (2018) However, in an epitaxially grown Mn3Sn/MgO(110) [001] bilayer structure, the energy landscape is found to possess two-fold symmetry with the emergence of a perpendicular magnetic anisotropy (PMA) owing to an epitaxial tensile strain. Higo et al. (2022); Yoon et al. (2023)
@@
Current-driven dynamics, including deterministic switching and chiral oscillations, have been experimentally explored in polycrystalline as well as epitaxial films of thicknesses ranging from sub-10 nm to 100 nm. Tsai et al. (2020); Takeuchi et al. (2021); Yan et al. (2022); Pal et al. (2022); Krishnaswamy et al. (2022); Higo et al. (2022); Yoon et al. (2023); Xu et al. (2023); Yoo et al. (2024); Zheng et al. (2025); Lee et al. (2025); Deng et al. (2023); Liu et al. (2023b) The AHE signal in the case of polycrystalline films, measured in a typical Hall bar setup, is considered to be an average of the final magnetic states in the various composing grains. Ikeda et al. (2018); Xie et al. (2022) While in some cases pure spin-orbit torque (SOT) is considered to be the fundamental mechanism behind the observed dynamics, Tsai et al. (2020); Takeuchi et al. (2021); Higo et al. (2022); Yoon et al. (2023); Deng et al. (2023) others have attributed the observed changes in the AHE signal to a combination of Joule heating and SOT (i.e., demagnetization of the polycrystalline Mn3Sn film as the temperature rises above $T_{\mathrm{N}}$ due to Joule heating, followed by SOT-driven re-magnetization as the film cools below $T_{\mathrm{N}}$ during the slow decrease in current. Pal et al. (2022); Krishnaswamy et al. (2022); Yoo et al. (2024) The SOT-driven oscillation Zhao, He, and Cai (2021); Shukla and Rakheja (2022); Lund et al. (2023) and switching Shukla, Qian, and Rakheja (2023); Xu et al. (2024); Shukla, Qian, and Rakheja (2024); Zhao et al. (2025); Deng et al. (2023) dynamics in both six-fold and two-fold energy degenerate Mn3Sn crystals have also been explored theoretically. In particular, these works have extended the ferromagnetic theory to shed light on the threshold current requirement for initiating dynamics as well as the dependence of the various threshold currents on the material properties. They have also explored the possible dependence of the oscillation frequency and the switching time on the input stimuli and the material parameters. The manipulation of the order parameter in Mn3Sn via SOT, together with a finite TMR in all-Mn3Sn junctions, can enable the development of AFM-based spintronic devices that are fully compatible with electronic circuitry.
@@
Figure 1: (a) Unit Cell of Mn3Sn with two Kagome planes. Mn spins in different planes are marked with different colors. (b) Octupole moment and relevant angles in Cartesian coordinate system.
@@
In this paper, we investigate the impact of thermal noise on the stability of a strained $\mathrm{Mn_{3}Sn}$ bit in the presence of an external magnetic field. We present a comprehensive analysis of the mechanisms underlying the free-energy symmetry breaking, examining the dependence of free energy on both the magnitude and orientation of the applied field, as well as the associated fluctuation modes. The interwell escape time of the magnetic octupole moment is theoretically evaluated as a function of the external field and temperature, with results corroborated by numerical simulations in the low-energy barrier regime ($\Delta E_{0}\leq{6}k_{\mathrm{B}}T$). We further extend our theoretical framework to encompass the relaxation dynamics of octupole moment and discuss prospective applications in random number generation and probabilistic computing.
Mn3Sn crystallizes into a hexagonal Kagome $D0_{19}$ lattice below its Néel temperature, exhibiting a slight excess of Mn atoms. Shukla, Qian, and Rakheja (2025) Sn atoms occupy the centers of the hexagons, whereas Mn atoms are located at the vertices. Each unit cell of Mn3Sn includes two Kagome planes with three Mn spins per plane. The magnetic moments of Mn atoms, forming an inverse triangular configuration, are slightly canted toward the in-plane easy axis, giving rise to a non-zero net magnetization as shown in Fig. 1(a). The magnetic ground state typically exhibits a six-fold degeneracy; Shukla, Qian, and Rakheja (2023) however, in the presence of an in-plane uniaxial strain, the symmetry is reduced to two-fold. Higo et al. (2022); Yoon et al. (2023) It is worth mentioning that intra-layer interactions are dominant as validated in Sec. I.2.3 of the Supplementary Material, so the following investigation focuses on one Kagome plane.
@@
$\mathcal{H}(\mathbf{m})$ $=J_{\mathrm{E}}((1+\delta_{\mathrm{E}})\mathbf{m}_{1}\cdot\mathbf{m}_{2}+\mathbf{m}_{2}\cdot\mathbf{m}_{3}+\mathbf{m}_{3}\cdot\mathbf{m}_{1})$ (1)
$+D_{\mathrm{M}}\mathbf{z}\cdot(\mathbf{m}_{1}\times\mathbf{m}_{2}+\mathbf{m}_{2}\times\mathbf{m}_{3}+\mathbf{m}_{3}\times\mathbf{m}_{1})$
$-\sum_{i=1}^{3}{\left[K_{\mathrm{u}}(\mathbf{m}_{i}\cdot\mathbf{u}_{i})^{2}+M_{\mathrm{s}}\mathbf{H}_{\mathrm{a}}\cdot\mathbf{m}_{i})\right]},$
@@
$\mathcal{H}(\theta_{\mathrm{oct}},\varphi_{\mathrm{oct}})=\frac{3}{4}M_{\mathrm{s}}H_{\mathrm{K}}\cos(2\varphi_{\mathrm{oct}})+\frac{3}{2}M_{\mathrm{s}}H_{\mathrm{J}}\cos^{2}\theta_{\mathrm{oct}}$ (2)
[…]
Here, we used $\mathbf{m}_{\mathrm{oct}}=\frac{1}{3}\mathcal{M}_{zx}\quantity[R\left(\frac{2\pi}{3}\right)\mathbf{m}_{1}+R\left(-\frac{2\pi}{3}\right)\mathbf{m}_{2}+\mathbf{m}_{3}]$ while $H_{\mathrm{J}}=\frac{(3J_{\mathrm{E}}+\sqrt{3}D_{\mathrm{M}})}{M_{s}}$ and $H_{\mathrm{K}}=-\frac{(4K_{\mathrm{u}}J_{\mathrm{E}}\delta_{\mathrm{E}})}{3M_{s}(J_{\mathrm{E}}+\sqrt{3}D_{\mathrm{M}})}$ are the strength of exchange field and uniaxial anisotropy field, respectively. $\theta_{\mathrm{H}}$ and $\varphi_{\mathrm{H}}$ are the polar and the azimuthal angle of the applied external field $\mathbf{H}_{\mathrm{a}}$, respectively. In addition, $\mathcal{M}_{z}x$ and $R$ represent mirror operation against $zx$ plane and anti-clockwise rotation operation against $z$ axis, respectively.
@@
$\delta_{\mathrm{E}}$ Strain parameter $2.6\times 10^{-4}$
@@
From Eq. (2), strong exchange interactions constrain $\mathbf{m}_{\mathrm{oct}}$ to lie within the Kagome plane, resulting in two degenerate energy maxima at $\varphi_{\mathrm{oct}}=0$ and $\pi$, and two minima at $\varphi_{\mathrm{oct}}=\pi/2$ and $3\pi/2$, in the absence of an external field. The equilibrium states near $\varphi_{\mathrm{oct}}=\pi/2$ and $3\pi/2$ are hereafter referred to as the up and down states, respectively. An applied external field breaks the symmetry of the local extrema. As shown in Fig. 2(a), when $\mathbf{H}_{\mathrm{a}}$ is applied along the x-x direction (i.e., $\varphi_{\mathrm{H}}=\pi$), the degeneracy of the energy maxima is lifted. For a film of volume $\mathcal{V}$, the energy barriers, $\Delta E_{\mathrm{x},1}$ and $\Delta E_{\mathrm{x},2}$, evaluated at $\varphi_{\mathrm{oct}}=0$ and $\pi$, respectively, are given as
@@
where $H_{\mathrm{x,eff}}=H_{\mathrm{ax}}\frac{(K_{\mathrm{u}}+J_{\mathrm{E}}\delta_{\mathrm{E}})}{3(J_{\mathrm{E}}+\sqrt{3}D_{\mathrm{M}})}$ is the normalized external field applied along the x-x direction with $H_{\mathrm{ax}}$ being the magnitude of the external field applied along x-x axis. In this configuration, the up and down states remain energetically equivalent; however, thermal fluctuations preferentially promote transitions across $\Delta E_{\mathrm{x,2}}$ due to its lower energy barrier. Furthermore, the two-state system reduces to one stable state at $\varphi_{\mathrm{oct}}=\pi$ as the magnitude of $H_{\mathrm{ax}}$ increases up to a certain magnitude, which for the parameters chosen here is 0.25 T.
In contrast, when the external field is applied along the y-y direction, it primarily perturbs the energy minima, as shown in Fig. 2(b). The energy maxima remain degenerate, resulting in symmetric energy barriers on either side of each state. The analytical expressions for the energy barriers $\Delta E_{\mathrm{y,1}}$ and $\Delta E_{\mathrm{y,2}}$, corresponding to the up and down states, respectively, are
@@
where $H_{\mathrm{y,eff}}=H_{\mathrm{ay}}\frac{(K_{\mathrm{u}}-J_{\mathrm{E}}\delta_{\mathrm{E}})}{3(J_{\mathrm{E}}+\sqrt{3}D_{\mathrm{M}})}$ denotes the normalized external field applied along the y-y direction with $H_{\mathrm{ay}}$ being the magnitude of the external field applied along y-y axis. When $H_{\mathrm{ay}}\geq 0.23~\mathrm{T}$, $\Delta E_{\mathrm{y,1}}$ vanishes and the two-state system reduces to one stable state at $\varphi_{\mathrm{oct}}=3\pi/2$. For comparison, Fig. 2(c) quantitatively illustrates the quadratic dependence of the energy barrier density on the magnitude of the external field. Additionally, we define $\Delta E_{0}$ as the energy barrier height in the absence of an external field, which serves as a reference in subsequent analyses.
Based on the harmonic transition-state theory (HTST), a a widely accepted framework for analyzing thermal behavior in ferromagnets Coffey and Kalmykov (2012); Morse and Feshbach (1946) and collinear antiferromagnets, Rózsa et al. (2019) the escape time with intermediate-to-high (IHD) damping is formulated per Eq. (5). In this equation, $\varepsilon_{j,\mathrm{min(sp)}}$ denotes the $j$-th eigenvalue of the Hessian matrix in the harmonic approximation near the energy minimum (saddle point), and $\lambda_{+}$ is the positive eigenvalue of the linearized Landau-Lifshitz-Gilbert (LLG) equation at the saddle point. $P_{\mathrm{min(sp)}}$ represents the number of Goldstone modes at the minimum (saddle point), and $V_{\mathrm{min(sp)}}$ denotes the corresponding phase-space volume. Additionally, $A(\delta E/k_{\mathrm{B}}T)$ is the depopulation factor that accounts for the coupling strength to the thermal bath, where $\delta E$ is the energy dissipated along an iso-energy contour encompassing the energy maximum. See Eq. (6) for $A(x)$.
In the theoretical analysis of escape time, we employ the angular variables $\theta_{\mathrm{oct}}$ and $\varphi_{\mathrm{oct}}$ to describe the dynamics of the octupole moment $\mathbf{m}_{\mathrm{oct}}$ within a perturbative framework. He and Liu (2024) The escape time due to thermal fluctuations over a single saddle point, from the up state to the down state under an external field $H_{\mathrm{ax}}$ or $H_{\mathrm{ay}}$ is given in Eq. (7), where $k=1$ and $k=2$ correspond to the magnetic octupole crossing the saddle point near $\varphi_{\mathrm{oct}}=0$ and $\pi$, respectively. The parameters $\beta_{x}$ and $\beta_{y}$ in Eq. (7) are given as $\left((-1)^{k}\frac{H_{\mathrm{x,eff}}}{H_{\mathrm{K}}}-1\right)$ and $\left(3\frac{H_{\mathrm{y,eff}}^{2}}{H_{\mathrm{K}}^{2}}-1\right)$, respectively. The derivation of Eqs. (5)-(7) is provided in Sec. I.1 of the Supplementary Material. Our calculations suggest that for the magnetic fields and energy barriers considered in this work, $\quantity[A(\delta E/k_{\mathrm{B}}T)]^{-1}$ can be approximated to unity with less than 1% error as shown in Sec. I.1.4 of the Supplementary Material.
@@
$A(x)=\exp\left(\frac{1}{2\pi}\int_{-\infty}^{\infty}\ln\left(1-e^{-x(0.25+y^{2})}\right)\frac{1}{0.25+y^{2}}\,dy\right).$ (6)
@@
Figure 3: The schematic of escape time from up state to down state while numerical extracted data is in markers and analytical predictions are plotted with lines. $\Delta E_{0}$ ranges from ${3}k_{\mathrm{B}}T$ to ${6}k_{\mathrm{B}}T$ and field is applied along (a) $x$-axis and (b) $y$-axis. (c) The time evolution of octupole moment with external field and $\Delta E_{0}=4k_{\mathrm{B}}T$.
@@
To validate Eqs. (7) and (8), the dynamics of the magnetic octupole is numerically explored via three coupled stochastic LLG equations, Mayergoyz, Bertotti, and Serpico (2009); Go et al. (2022) for $\Delta E_{0}$ ranging from $k_{\mathrm{B}}T$ to $5k_{\mathrm{B}}T$ and $H_{\mathrm{a}}\leq 0.1~\mathrm{T}$. For each sublattice, $i$ (= 1,2,3), the LLG equation is given as $\partialderivative{\mathbf{m}_{i}}{t}=-\gamma(\mathbf{m}_{i}\crossproduct\mathbf{H}_{i}^{\mathrm{eff}})+\alpha\quantity(\mathbf{m}_{i}\times\partialderivative{\mathbf{m}_{i}}{t})$, where $\mathbf{H}_{i}^{\mathrm{eff}}=-\frac{1}{M_{\mathrm{s}}}\frac{\partial\mathcal{H}}{\partial\mathbf{m}_{i}}+\mathbf{H}_{\mathrm{thermal},i}$ is the effective magnetic field and includes the thermal noise field, $\mathbf{H}_{\mathrm{thermal},i}=\sqrt{\frac{2\alpha k_{\mathrm{B}}T}{\gamma\mu_{0}M_{\mathrm{s}}^{2}\mathcal{V}\Delta t}}\bm{\xi}^{i}_{t}$. In this work, we assume $\mathbf{H}_{\mathrm{thermal},i}$ to be Gaussian with zero mean as well as uncorrelated in space, time, and between sublattices, such that $\bm{\xi}^{i}_{t}\sim\mathcal{N}(0,1)$Go et al. (2022); Ament et al. (2016) Here, $\gamma$ and $\mu_{0}$ denote the gyromagnetic ratio and vacuum permeability, respectively, while $\Delta t$ is the time step used in the simulation. Other parameters are defined in Table 1. The coupled LLG equations are solved using the finite difference method implemented within a self-developed CUDA framework, which has been benchmarked against published work with Konakanchi et al. (2025) and without Shukla, Qian, and Rakheja (2024) thermal noise. Additional details of the numerical LLG framework, its validation, and numerical extraction of escape time are provided in Sec. I.2 of the Supplementary Material.